# Improved 3D Scene Stylization via Text-Guided Generative Image Editing with Region-Based Control

### v1

@@ Abstract
Recent advances in text-driven 3D scene editing and stylization, which leverage the powerful capabilities of 2D generative models, have demonstrated promising outcomes. However, challenges remain in ensuring high-quality stylization and view consistency simultaneously. Moreover, applying style consistently to different regions or objects in the scene with semantic correspondence is a challenging task. To address these limitations, we introduce techniques that enhance the quality of 3D stylization while maintaining view consistency and providing optional region-controlled style transfer. Our method achieves stylization by re-training an initial 3D representation using stylized multi-view 2D images of the source views. Therefore, ensuring both style consistency and view consistency of stylized multi-view images is crucial. We achieve this by extending the style-aligned depth-conditioned view generation framework, replacing the fully shared attention mechanism with a single reference-based attention-sharing mechanism, which effectively aligns style across different viewpoints. Additionally, inspired by recent 3D inpainting methods, we utilize a grid of multiple depth maps as a single-image reference to further strengthen view consistency among stylized images. Finally, we propose Multi-Region Importance-Weighted Sliced Wasserstein Distance Loss, allowing styles to be applied to distinct image regions using segmentation masks from off-the-shelf models. We demonstrate that this optional feature enhances the faithfulness of style transfer and enables the mixing of different styles across distinct regions of the scene. Experimental evaluations, both qualitative and quantitative, demonstrate that our pipeline effectively improves the results of text-driven 3D stylization. {CCSXML} <ccs2012> <concept> <concept_id>10010147.10010371.10010372.10010375</concept_id> <concept_desc>Computing methodologies Non-photorealistic rendering</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10010147.10010178.10010224.10010240</concept_id> <concept_desc>Computing methodologies Computer vision representations</concept_desc> <concept_significance>500</concept_significance> </concept> </ccs2012>
@@ 1 Introduction
Recent 3D reconstruction techniques, such as Neural Radiance Field (NeRF) [MST20] and Gaussian splatting [KKLD23], enable the creation of high-quality digital representations of real scenes from 2D photos. Editing and stylizing such 3D representations have emerged as an important research focus for applications such as content creation and 3D post-editing. However, direct manipulation of 3D representations is not straightforward as multi-view consistency of both geometry and color need to considered. To alleviate such a bottleneck and bypass the necessity for scarce 3D data, recent efforts often utilize foundational 2D generative models such as Stable Diffusion to achieve 3D editing and stylization via knowledge distillation. While these methods have pioneered 3D editing, most techniques rely on iterative techniques such as SDS [PJBM22] and iterative dataset updates [HTE23, VH24] which often require significant time until convergence. As an alternative, some techniques [FMH24, CLV24] propose a more simplified sequential process: (a) consistent multi-view edit, and then (b) 3D re-train on edited images. This approach benefits users by providing a preview of the stylized scene before full 3D re-training.
In our work, we also follow a generate-then-train approach similar to Style-NeRF2NeRF [FMH24] where we first stylize 2D multi-view images and then fine-tune the source Gaussian Splatting [KKLD23, HYC24] scene based on these images. We observe that consistency of stylized multi-view images contributes to effective 3D stylization. To this end, we propose a training-free technique that uses a text-driven generation pipeline conditioned on tiled depth maps for consistent view stylization.
Along with the above improvements, we propose a multi-region importance-weighted sliced Wasserstein distance loss (MR-IW-SWD) to perform semantically consistent style transfer. While prior work [FMH24] effectively applies SWD for 3D style transfer, it lacks region-specific control—a key limitation in NeRF-based representations, where stylization is applied uniformly across sampled patches. To address this, we extend the SWD loss with segmentation masks for semantically aware, region-specific style transfer. Additionally, we adopt importance-weighted sampling to improve efficiency by prioritizing perceptually salient projection directions. These enhancements, integrated into our MR-IW-SWD, enable efficient and semantically consistent 3D stylization. Our contributions are summarized as following:
@@ 2 Related Work
Diffusion models [SDWMG15, SSDK20, DN21] are powerful generative models recognized for producing high-quality, diverse images. Drawing on concepts from non-equilibrium thermodynamics, these models generate images by reversing a diffusion process, gradually denoising from random noise to a coherent image. Typically, they employ classifier-free guidance [HS22] for text-conditioned image generation. Extensions like ControlNet [ZRA23] allow for more detailed control by incorporating additional conditions, such as depth maps, normal maps, and feature lines. Similarly, IP-Adapter [YZL23] enables enhanced control by using an image as a prompt.
@@ 2.2 Style Transfer
Style transfer blends a source image with a style image, creating a result that retains the content of the source but reflects the style of the other. Since the foundational neural algorithm by Gatys et al. [GEB15], research has focused on enhancing 2D style transfer including faster optimization [JAFF16, HB17], zero-shot transfer [LFY17], and photorealism [LPSB17]. While most previous methods leveraged the feature distributions of a pretrained convolutional model such as VGG19 [SZ14] to manipulate style, recent advances in diffusion models have introduced enhanced control over stylized image generation using both text and reference images. Lightweight fine-tuning methods [FVSCO24, SRC23, EBA23, SJB24, KZL23] adapt the underlying diffusion models for personalized stylization. Pre-trained add-ons [XWS24, YZL23] allow flexible style manipulation by referencing text or an image directly. Training-free approaches based on latent code inversion [MHA23, CLPL23, TGBD23, PKSZ23] also facilitate style control, while others [CWQ23, HVFCO24, HMT22, HACO23, MHS21] use attention modulation for controlled generation without finetuning. Our work leverages a training-free approach with attention-sharing [HVFCO24] to generate consistently stylized images across multi-views.
@@
Modern 3D representations such as Neural Radiance Fields (NeRF) [MST20] and 3D Gaussian Splatting [KKLD23] have become increasingly popular due to their advantages over traditional approaches like polygon meshes or voxel grids. NeRF, as an implicit representation, encodes scenes in a continuous volumetric function, enabling photo-realistic novel view synthesis at arbitrary resolutions with a compact memory footprint. More recently, 3D Gaussian Splatting has emerged as a powerful explicit alternative that supports real-time, high-quality rendering by leveraging conventional graphics pipelines. In this work, we adopt 2D Gaussian Splatting [HYC24] as our base representation—a recent variant that achieves superior geometry reconstruction while maintaining competitive performance and view synthesis quality.
@@
Recent research has extended 2D style transfer techniques to 3D by embedding deep feature statistics into volumetric scene representations, particularly with NeRF-based models [LZC23, WJC23, ZKB22, CTT22, HHY22, NPLX22, PHY23, KYO24, ZHX23] and more recently with Gaussian Splatting [GLC24, CCZ24, WFZ24, CLV24, WYW24, LZX24, GWRM25, KHR24]. Beyond reference image-based stylization, text-driven 3D editing methods have emerged, leveraging the generative power of 2D text-to-image (T2I) models. Some methods [HTE23, DW24] alternate between scene finetuning and multi-view supervision, while others such as [KPS24] directly update the underlying NeRF representation in a single optimization loop. In contrast, generate-then-train approaches [DEL24, FMH24, CLV24, WQP25] decouple the process by first generating stylized views and then refining the 3D scene. This strategy offers a more intuitive and interactive workflow by allowing users to preview stylization effects before full 3D refinement. In our work, we adopt this sequential paradigm and propose a training-free diffusion pipeline for generating style-consistent multi-view images. These images are then used to fine-tune the 3D scene using our improved style transfer loss, resulting in high-quality and semantically coherent 3D stylization.
@@ 3.1.1 Gaussian Splatting
3D Gaussian Splatting (3DGS) represents a scene using a set of anisotropic 3D Gaussians, which are typically initialized from a sparse point cloud reconstructed via Structure-from-Motion (SfM) [SF16]. Each Gaussian is explicitly parameterized by its 3D position $\mathbf{x}_{k}$ and a covariance matrix $\Sigma$ defined as $\Sigma=\mathbf{R}\mathbf{S}\mathbf{S}^{\top}\mathbf{R}^{\top}$:
@@
where $\mathbf{S}$ is a scaling matrix and $\mathbf{R}$ is a rotation matrix. To render an image, the Gaussians are transformed into camera coordinates using the world-to-camera transformation matrix $\mathbf{W}$ and projected onto the image plane via a local affine projection $\mathbf{J}$ [ZPVBG01], yielding a transformed covariance $\Sigma^{\prime}=\mathbf{J}\mathbf{W}\Sigma\mathbf{W}^{\top}\mathbf{J}^{\top}$. In this work, we adopt 2D Gaussian Splatting (2DGS) [HB17] as our base representation, which is a variant that offers view-consistent geometry. Specifically, 2DGS omits the third row and column of $\Sigma^{\prime}$, and represents the projected Gaussians $g^{2D}$ using a 2D covariance matrix $\Sigma^{2D}$. Rendering is achieved via volumetric alpha blending, integrating view-dependent appearance values $\mathbf{c}_{k}$ from front to back, weighted by their corresponding alpha values $\alpha_{k}$, as follows:
@@ 3.1.2 Multi-View Stylization
Recent works have explored leveraging powerful 2D generative models, such as diffusion models, for text-driven 3D scene editing and stylization. While these methods show promising results, challenges persist in ensuring high-quality stylization, maintaining view consistency, and applying style consistently across different regions or objects within a scene. Style-NeRF2NeRF [FMH24] proposes an effective pipeline to address these challenges by stylizing a 3D scene using 2D image diffusion models [PEL23]. The pipeline begins with a Neural Radiance Field (NeRF) model reconstructed from a set of multi-view images. Style transfer is then achieved by refining the source NeRF model using stylized images generated by a style-aligned image-to-image diffusion model. A key step involves generating perceptually similar multi-view images based on a given target style prompt. This is accomplished by employing a depth-conditioned diffusion model enhanced with an attention-sharing mechanism. Following the generation of stylized multi-view images, the style transfer process is guided using the sliced Wasserstein loss, which is computed based on feature maps extracted from a pre-trained Convolutional Neural Network (CNN) model [SZ14]. A notable aspect of Style-NeRF2NeRF pipeline is its decoupled nature, allowing users the flexibility to experiment with various style prompts and preview the stylized 3D results before proceeding to the NeRF fine-tuning stage. This method has demonstrated the ability to transfer diverse artistic styles to real-world 3D scenes while achieving competitive quality.
@@ 3.1.3 Sliced Wasserstein Distance-Based Style Transfer
Feature statistics from pre-trained Convolutional Neural Networks (CNNs), such as VGG-19 [SZ14], are well-known for capturing the style of an image [GEB15, JAFF16, HB17, LFY17, LPSB17]. Sliced Wasserstein Distance (SWD) loss has been proposed as an effective alternative for learning image style, especially in applications such as texture synthesis [HVCB21] and 3D style transfer [FMH24]. In our work, we build upon SWD to lift the stylized 2D multi-view images into a 3D representation with additional spatial control.
@@ 3.2 Multi-View Editing With Tiled Depth Reference
Modern image generation models [RBL22, PEL23, EKB24] exhibit a degree of 3D awareness, attributed to training on large-scale image-text datasets [ZZXZ23b, ZZXZ23a]. Conditioning these models with 3D cues such as depth maps via ControlNet [ZRA23] has been shown to significantly improve multi-view consistency. Building on this capability, several recent methods [WHJ24, DEL24] employ inpainting pipelines over grids of image priors to achieve view-consistent edits. Inspired by this strategy, we instead construct a tiled grid of sampled depth maps $D_{ref}$ as a unified reference input for our depth-conditioned multi-view generation pipeline. Similar to Style-NeRF2NeRF [FMH24], we incorporate an attention-sharing mechanism [HVFCO24] to propagate consistent style across views. Unlike the fully-shared attention approach proposed in prior work, however, our method anchors the generation process on tiled depth maps representing multiple views (arranged two-by-two in our experiments) representing different views, enforcing structural alignment across diverse camera angles. Empirically, we find that this reference-guided attention sharing further enhances appearance consistency, outperforming the closest prior method [FMH24].
Following [HVFCO24], we replace the self-attention operation in the T2I diffusion model using the following formulas:
@@
where target queries $\hat{Q}_{t}$ and keys $\hat{K}_{t}$ are normalized via adaptive instance normalization [HB17] with reference queries $Q_{ref}$ and keys $K_{ref}$ of the tiled control reference $D_{ref}$ :

### v2

@@ Abstract
Recent advances in text-driven 3D scene editing and stylization, which leverage the powerful capabilities of 2D generative models, have demonstrated promising outcomes. However, challenges remain in ensuring high-quality stylization and view consistency simultaneously. Moreover, applying style consistently to different regions or objects in the scene with semantic correspondence is a challenging task. To address these limitations, we introduce techniques that enhance the quality of 3D stylization while maintaining view consistency and providing optional region-controlled style transfer. Our method achieves stylization by re-training an initial 3D representation using stylized multi-view 2D images of the source views. Therefore, ensuring both style consistency and view consistency of stylized multi-view images is crucial. We achieve this by extending the style-aligned depth-conditioned view generation framework, replacing the fully shared attention mechanism with a single reference-based attention-sharing mechanism, which effectively aligns style across different viewpoints. Additionally, inspired by recent 3D inpainting methods, we utilize a grid of multiple depth maps as a single-image reference to further strengthen view consistency among stylized images. Finally, we propose Multi-Region Importance-Weighted Sliced Wasserstein Distance Loss, allowing styles to be applied to distinct image regions using segmentation masks from off-the-shelf models. We demonstrate that this optional feature enhances the faithfulness of style transfer and enables the mixing of different styles across distinct regions of the scene. Experimental evaluations, both qualitative and quantitative, demonstrate that our pipeline effectively improves the results of text-driven 3D stylization. Project page {CCSXML} <ccs2012> <concept> <concept_id>10010147.10010371.10010372.10010375</concept_id> <concept_desc>Computing methodologies Non-photorealistic rendering</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10010147.10010178.10010224.10010240</concept_id> <concept_desc>Computing methodologies Computer vision representations</concept_desc> <concept_significance>500</concept_significance> </concept> </ccs2012>
@@ 1 Introduction
Recent 3D reconstruction techniques, such as Neural Radiance Field (NeRF) [mildenhall2020nerf] and Gaussian splatting [kerbl3Dgaussians], enable the creation of high-quality digital representations of real scenes from 2D photos. Editing and stylizing such 3D representations have emerged as an important research focus for applications such as content creation and 3D post-editing. However, direct manipulation of 3D representations is not straightforward as multi-view consistency of both geometry and color need to considered. To alleviate such a bottleneck and bypass the necessity for scarce 3D data, recent efforts often utilize foundational 2D generative models such as Stable Diffusion to achieve 3D editing and stylization via knowledge distillation. While these methods have pioneered 3D editing, most techniques rely on iterative techniques such as SDS [poole2022dreamfusion] and iterative dataset updates [instructnerf2023, igs2gs] which often require significant time until convergence. As an alternative, some techniques [fujiwara2024sn2n, chen2024dge] propose a more simplified sequential process: (a) consistent multi-view edit, and then (b) 3D re-train on edited images. This approach benefits users by providing a preview of the stylized scene before full 3D re-training.
In our work, we also follow a generate-then-train approach similar to Style-NeRF2NeRF [fujiwara2024sn2n] where we first stylize 2D multi-view images and then fine-tune the source Gaussian Splatting [kerbl3Dgaussians, huang20242d] scene based on these images. We observe that consistency of stylized multi-view images contributes to effective 3D stylization. To this end, we propose a training-free technique that uses a text-driven generation pipeline conditioned on tiled depth maps for consistent view stylization.
Along with the above improvements, we propose a multi-region importance-weighted sliced Wasserstein distance loss (MR-IW-SWD) to perform semantically consistent style transfer. While prior work [fujiwara2024sn2n] effectively applies SWD for 3D style transfer, it lacks region-specific control—a key limitation in NeRF-based representations, where stylization is applied uniformly across sampled patches. To address this, we extend the SWD loss with segmentation masks for semantically aware, region-specific style transfer. Additionally, we adopt importance-weighted sampling to improve efficiency by prioritizing perceptually salient projection directions. These enhancements, integrated into our MR-IW-SWD, enable efficient and semantically consistent 3D stylization. Our contributions are summarized as following:
@@ 2 Related Work
Diffusion models [sohl2015deep, song2020score, dhariwal2021diffusion] are powerful generative models recognized for producing high-quality, diverse images. Drawing on concepts from non-equilibrium thermodynamics, these models generate images by reversing a diffusion process, gradually denoising from random noise to a coherent image. Typically, they employ classifier-free guidance [ho2022classifier] for text-conditioned image generation. Extensions like ControlNet [zhang2023adding] allow for more detailed control by incorporating additional conditions, such as depth maps, normal maps, and feature lines. Similarly, IP-Adapter [ye2023ip-adapter] enables enhanced control by using an image as a prompt.
@@ 2.2 Style Transfer
Style transfer blends a source image with a style image, creating a result that retains the content of the source but reflects the style of the other. Since the foundational neural algorithm by Gatys et al. [gatys2015neural], research has focused on enhancing 2D style transfer including faster optimization [johnson2016perceptual_faststyle, huang2017arbitrary_adain], zero-shot transfer [li2017universal], and photorealism [luan2017deepphoto]. While most previous methods leveraged the feature distributions of a pretrained convolutional model such as VGG19 [simonyan2014very] to manipulate style, recent advances in diffusion models have introduced enhanced control over stylized image generation using both text and reference images. Lightweight fine-tuning methods [frenkel2024implicit, shah2023ziplora, Everaert_2023_ICCV, sohn2024styledrop, kawar2023imagic] adapt the underlying diffusion models for personalized stylization. Pre-trained add-ons [xing2024csgo, ye2023ip-adapter] allow flexible style manipulation by referencing text or an image directly. Training-free approaches based on latent code inversion [mokady2023null, cheng2023general_i2i, tumanyan2023plug, parmar2023zero_i2i] also facilitate style control, while others [cao2023masactrl, hertz2024style, hertz2022prompt, hertz2023delta, meng2021sdedit] use attention modulation for controlled generation without finetuning. Our work leverages a training-free approach with attention-sharing [hertz2024style] to generate consistently stylized images across multi-views.
@@
Modern 3D representations such as Neural Radiance Fields (NeRF) [mildenhall2020nerf] and 3D Gaussian Splatting [kerbl3Dgaussians] have become increasingly popular due to their advantages over traditional approaches like polygon meshes or voxel grids. NeRF, as an implicit representation, encodes scenes in a continuous volumetric function, enabling photo-realistic novel view synthesis at arbitrary resolutions with a compact memory footprint. More recently, 3D Gaussian Splatting has emerged as a powerful explicit alternative that supports real-time, high-quality rendering by leveraging conventional graphics pipelines. In this work, we adopt 2D Gaussian Splatting [huang20242d] as our base representation—a recent variant that achieves superior geometry reconstruction while maintaining competitive performance and view synthesis quality.
@@
Recent research has extended 2D style transfer techniques to 3D by embedding deep feature statistics into volumetric scene representations, particularly with NeRF-based models [liu2023stylerf, wang2023nerfart, zhang2022arf, chiang2022stylizing, huang2022stylizednerf, nguyen2022snerf, pang2023locally, kim2024fprf, zhang2023ref] and more recently with Gaussian Splatting [gu2024artnvg, chen2024gaussianeditor, wang2024gaussianeditor, chen2024dge, wang2024view, liu2024stylegaussian, galerne2025sgsst, kovacs2024]. Beyond reference image-based stylization, text-driven 3D editing methods have emerged, leveraging the generative power of 2D text-to-image (T2I) models. Some methods [instructnerf2023, dong2024vica] alternate between scene finetuning and multi-view supervision, while others such as [Koo:2024PDS] directly update the underlying NeRF representation in a single optimization loop. In contrast, generate-then-train approaches [signerf, fujiwara2024sn2n, chen2024dge, wynn2025morpheus] decouple the process by first generating stylized views and then refining the 3D scene. This strategy offers a more intuitive and interactive workflow by allowing users to preview stylization effects before full 3D refinement. In our work, we adopt this sequential paradigm and propose a training-free diffusion pipeline for generating style-consistent multi-view images. These images are then used to fine-tune the 3D scene using our improved style transfer loss, resulting in high-quality and semantically coherent 3D stylization.
@@ 3.1.1 Gaussian Splatting
3D Gaussian Splatting (3DGS) represents a scene using a set of anisotropic 3D Gaussians, which are typically initialized from a sparse point cloud reconstructed via Structure-from-Motion (SfM) [schoenberger2016sfm]. Each Gaussian is explicitly parameterized by its 3D position $\mathbf{x}_{k}$ and a covariance matrix $\Sigma$ defined as $\Sigma=\mathbf{R}\mathbf{S}\mathbf{S}^{\top}\mathbf{R}^{\top}$:
@@
where $\mathbf{S}$ is a scaling matrix and $\mathbf{R}$ is a rotation matrix. To render an image, the Gaussians are transformed into camera coordinates using the world-to-camera transformation matrix $\mathbf{W}$ and projected onto the image plane via a local affine projection $\mathbf{J}$ [zwicker2001ewa], yielding a transformed covariance $\Sigma^{\prime}=\mathbf{J}\mathbf{W}\Sigma\mathbf{W}^{\top}\mathbf{J}^{\top}$. In this work, we adopt 2D Gaussian Splatting (2DGS) [huang2017arbitrary_adain] as our base representation, which is a variant that offers view-consistent geometry. Specifically, 2DGS omits the third row and column of $\Sigma^{\prime}$, and represents the projected Gaussians $g^{2D}$ using a 2D covariance matrix $\Sigma^{2D}$. Rendering is achieved via volumetric alpha blending, integrating view-dependent appearance values $\mathbf{c}_{k}$ from front to back, weighted by their corresponding alpha values $\alpha_{k}$, as follows:
@@ 3.1.2 Multi-View Stylization
Recent works have explored leveraging powerful 2D generative models, such as diffusion models, for text-driven 3D scene editing and stylization. While these methods show promising results, challenges persist in ensuring high-quality stylization, maintaining view consistency, and applying style consistently across different regions or objects within a scene. Style-NeRF2NeRF [fujiwara2024sn2n] proposes an effective pipeline to address these challenges by stylizing a 3D scene using 2D image diffusion models [podell2023sdxl]. The pipeline begins with a Neural Radiance Field (NeRF) model reconstructed from a set of multi-view images. Style transfer is then achieved by refining the source NeRF model using stylized images generated by a style-aligned image-to-image diffusion model. A key step involves generating perceptually similar multi-view images based on a given target style prompt. This is accomplished by employing a depth-conditioned diffusion model enhanced with an attention-sharing mechanism. Following the generation of stylized multi-view images, the style transfer process is guided using the sliced Wasserstein loss, which is computed based on feature maps extracted from a pre-trained Convolutional Neural Network (CNN) model [simonyan2014very]. A notable aspect of Style-NeRF2NeRF pipeline is its decoupled nature, allowing users the flexibility to experiment with various style prompts and preview the stylized 3D results before proceeding to the NeRF fine-tuning stage. This method has demonstrated the ability to transfer diverse artistic styles to real-world 3D scenes while achieving competitive quality.
@@ 3.1.3 Sliced Wasserstein Distance-Based Style Transfer
Feature statistics from pre-trained Convolutional Neural Networks (CNNs), such as VGG-19 [simonyan2014very], are well-known for capturing the style of an image [gatys2015neural, johnson2016perceptual_faststyle, huang2017arbitrary_adain, li2017universal, luan2017deepphoto]. Sliced Wasserstein Distance (SWD) loss has been proposed as an effective alternative for learning image style, especially in applications such as texture synthesis [heitz2021sliced] and 3D style transfer [fujiwara2024sn2n]. In our work, we build upon SWD to lift the stylized 2D multi-view images into a 3D representation with additional spatial control.
@@ 3.2 Multi-View Editing With Tiled Depth Reference
Modern image generation models [rombach2022high, podell2023sdxl, esser2024scaling] exhibit a degree of 3D awareness, attributed to training on large-scale image-text datasets [zhan2023does, zhan2023general]. Conditioning these models with 3D cues such as depth maps via ControlNet [zhang2023adding] has been shown to significantly improve multi-view consistency. Building on this capability, several recent methods [weber2024nerfiller, signerf] employ inpainting pipelines over grids of image priors to achieve view-consistent edits. Inspired by this strategy, we instead construct a tiled grid of sampled depth maps $D_{ref}$ as a unified reference input for our depth-conditioned multi-view generation pipeline. Similar to Style-NeRF2NeRF [fujiwara2024sn2n], we incorporate an attention-sharing mechanism [hertz2024style] to propagate consistent style across views. Unlike the fully-shared attention approach proposed in prior work, however, our method anchors the generation process on tiled depth maps representing multiple views (arranged two-by-two in our experiments) representing different views, enforcing structural alignment across diverse camera angles. Empirically, we find that this reference-guided attention sharing further enhances appearance consistency, outperforming the closest prior method [fujiwara2024sn2n].
Following [hertz2024style], we replace the self-attention operation in the T2I diffusion model using the following formulas:
@@
where target queries $\hat{Q}_{t}$ and keys $\hat{K}_{t}$ are normalized via adaptive instance normalization [huang2017arbitrary_adain] with reference queries $Q_{ref}$ and keys $K_{ref}$ of the tiled control reference $D_{ref}$ :